OpenStudy (anonymous):

Solve the equation 10-5b= -8 -14b 2/19 18/19 -2/19 -2

OpenStudy (hoblos):

10-5b= -8 -14b add 14b 10+ 9b = -8 subtract 10 9b = -18 divide by 9 b=-2